OpenStudy (anonymous):

find the number of each sentence follow the order of operations. 54+24(divide) 3-3-30=

OpenStudy (anonymous):

i think it's probably 9000

OpenStudy (anonymous):

or maybe over

OpenStudy (anonymous):

\[\frac{ 54 + 24 }{ 3-3-30 }\]

OpenStudy (anonymous):

=-2.6

OpenStudy (anonymous):

You will add first so 54 + 24 = 78 ---- 3 - 3 - 30 Now subtract 3-3 = 0 then 0 - 30 = -30 78 / -30 = 2.6

OpenStudy (anonymous):

Thats what microsoft 64bit calculator says, so it is inevitably correct!

OpenStudy (anonymous):

I mean -2.6

OpenStudy (anonymous):

Its really simple even without calculator. Just use equations.

OpenStudy (anonymous):

Hope that helped you 'understand' what to do :)